What is the positive root of the equation 3x²-27=0?
1)9
2)-9
3)3
4)-3

Answer by anantha(86)

sol:
3x^2-27=0
adding 27 on both sides
3x^2-27+27=0+27
3x^2=27
dividing by 3 on both sides
3x^2/3=27/3
x^2=9
x=sqrt(9) (i.e;sqrt(9)=sqrt(3*3)=3)
x=plus or minus 3
the positive root of given equation is 3